OpenStudy (anonymous):

George collected $48 for selling 5 sweatshirts and 2 T-shirts. Toni collected $32 for selling 3 sweatshirts and 2 T-shirts. How do I put this in Substitution method?

OpenStudy (anonymous):

What you can do is make it into one equation. $80=8 sweatshirts + 4 t-shirts (so in total 12 items were sold) 8/12=75% 4/12=25% (.75)x80= 60 (.25)x80=20 60=8s (s=sweatshirts) 20=4t (t=T-shirts) 60/8=$7.5 for sweatshirts 20/4=$5 for T-shirts

OpenStudy (anonymous):

that makes no since what so ever

OpenStudy (anonymous):

i need two equations not one

OpenStudy (anonymous):

Hold on, I'll go do it on paper then come back to you